# ULog: Unsupervised Log Parsing with Large Language Models through Log Contrastive Units

### v1

@@ 6. THREATS TO VALIDITY
Data leakage Given that large language models (LLMs) are trained on extensive datasets, one potential risk is data leakage. Specifically, the LLM used in ULog might have been trained on open-source log datasets, which could result in memorizing ground-truth templates rather than performing genuine inference. However, experimental results indicate that the performance of merely using LLMs (LILAC w/o ICL) is significantly lower compared to ULog, suggesting a low likelihood of direct memorization. Additionally, ULog utilizes the same LLM, i.e., gpt-turbo-3.5-0613 model for experiments, aligning with previous work, LILAC.
Randomness Randomness can influence the performance of ULog and other baseline methods. To mitigate this issue, we minimized the randomness of the LLM by setting the temperature to 0, ensuring consistent outputs for the same input text. Additionally, we conducted each experiment five times for every experimental setting and used the average of these results as the final outcome.
Implementation and settings To mitigate the bias of implementation and settings, in our evaluation, we compared our ULog with state-of-the-art approaches within the same evaluation framework. We adopted the implementations from their replication packages and benchmarks, using the parameters and settings (e.g., number of log templates and similarity threshold) optimized by previous work (Zhu et al., 2019; Jiang et al., 2023b). Moreover, the results of the baseline approaches are consistent with the best results in recent benchmarks.
@@ 7. Related Works
On the other hand, semantic-based log parsers utilize neural networks (Li et al., 2023a; Liu et al., 2022) or language models (Le and Zhang, 2023, 2022) to identify log templates and parameters by understanding the semantics of log messages. They require human-labelled log templates to train or tune the models by learning the semantics and patterns in the labelled log data. UniParser (Liu et al., 2022) is one of the pioneering work in parsing logs with a focus on their semantic meaning. It integrates a BiLSTM-based semantics miner with a joint parser to identify log templates. Additionally, LogPPT (Le and Zhang, 2023) proposed identifying log templates and parameters using prompt-based few-shot learning, based on the RoBERTa model. Recently, with the rise of large language models (LLMs), a series of LLM-based log parsers have been developed to achieve more effective log parsing. These LLM-based log parsers leverage fine-tuning (Liu et al., 2023b) or in-context learning (Xu et al., 2024; Jiang et al., 2023a) to specialize LLMs for log parsing tasks, thereby achieving remarkable performance. However, these semantic-based log parsers heavily rely on labelled data, which limits their ability to generalize to different types of log data or evolving log data (Jiang et al., 2023b). In contrast, our proposed unsupervised LLM-based log parser, ULog, does not require labelled examples, allowing it to generalize to diverse and evolving log data.
@@ 8. Conclusion
In this work, we propose an LLM-based unsupervised log parser named ULog, which leverages log contrastive units (LCUs) to facilitate effective comparisons by the LLM. To efficiently identify effective LCUs from large-scale log data, ULog employs a hierarchical sharder to divide logs into buckets, thereby reducing sampling overhead and enabling parallel computation. Additionally, ULog incorporates a hybrid LCU selector that jointly measures the commonality and variability of LCUs, which is crucial for prompting LLMs. Furthermore, ULog introduces an improved prompt format to guide LLMs in a zero-shot setting. Experimental results on 14 large-scale public datasets demonstrate that ULog achieves high parsing accuracy, significantly outperforming other unsupervised parsers and being comparable to state-of-the-art parsers that require labelled data. In terms of efficiency, ULog is also comparable to the fastest baseline parser, highlighting its potential for application in real-world systems.

### v2

@@ 6. THREATS TO VALIDITY
Data leakage Given that large language models (LLMs) are trained on extensive datasets, one potential risk is data leakage. Specifically, the LLM used in LUNAR might have been trained on open-source log datasets, which could result in memorizing ground-truth templates rather than performing genuine inference. However, experimental results indicate that the performance of merely using LLMs (LILAC w/o ICL) is significantly lower compared to LUNAR, suggesting a low likelihood of direct memorization. Additionally, LUNAR utilizes the same LLM, i.e., gpt-turbo-3.5-0613 model for experiments, aligning with previous work, LILAC.
Randomness Randomness can influence the performance of LUNAR and other baseline methods. To mitigate this issue, we minimized the randomness of the LLM by setting the temperature to 0, ensuring consistent outputs for the same input text. Additionally, we conducted each experiment five times for every experimental setting and used the average of these results as the final outcome.
Implementation and settings To mitigate the bias of implementation and settings, in our evaluation, we compared our LUNAR with state-of-the-art approaches within the same evaluation framework. We adopted the implementations from their replication packages and benchmarks, using the parameters and settings (e.g., number of log templates and similarity threshold) optimized by previous work (Zhu et al., 2019; Jiang et al., 2023b). Moreover, the results of the baseline approaches are consistent with the best results in recent benchmarks.
@@ 7. Related Works
On the other hand, semantic-based log parsers utilize neural networks (Li et al., 2023b; Liu et al., 2022) or language models (Le and Zhang, 2023) to identify log templates and parameters by understanding the semantics of log messages. They require human-labelled log templates to train or tune the models by learning the semantics and patterns in the labelled log data. UniParser (Liu et al., 2022) is one of the pioneering work in parsing logs with a focus on their semantic meaning. It integrates a BiLSTM-based semantics miner with a joint parser to identify log templates. Additionally, LogPPT (Le and Zhang, 2023) proposed identifying log templates and parameters using prompt-based few-shot learning, based on the RoBERTa model. Recently, with the rise of large language models (LLMs), a series of LLM-based log parsers have been developed to achieve more effective log parsing. These LLM-based log parsers leverage fine-tuning (Liu et al., 2023b) or in-context learning (Xu et al., 2024; Jiang et al., 2023a) to specialize LLMs for log parsing tasks, thereby achieving remarkable performance. However, these semantic-based log parsers heavily rely on labelled data, which limits their ability to generalize to different types of log data or evolving log data (Jiang et al., 2023b). In contrast, our proposed unsupervised LLM-based log parser, LUNAR, does not require labelled examples, allowing it to generalize to diverse and evolving log data.
@@ 8. Conclusion
In this work, we propose an LLM-based unsupervised log parser named LUNAR, which leverages log contrastive units (LCUs) to facilitate effective comparisons by the LLM. To efficiently identify effective LCUs from large-scale log data, LUNAR employs a hierarchical sharder to divide logs into buckets, thereby reducing sampling overhead and enabling parallel computation. Additionally, LUNAR incorporates a hybrid LCU selector that jointly measures the commonality and variability of LCUs, which is crucial for prompting LLMs. Furthermore, LUNAR introduces an improved prompt format to guide LLMs in a zero-shot setting. Experimental results on 14 large-scale public datasets demonstrate that LUNAR achieves high parsing accuracy, significantly outperforming other unsupervised parsers and being comparable to state-of-the-art parsers that require labelled data. In terms of efficiency, LUNAR is also comparable to the fastest baseline parser, highlighting its potential for application in real-world systems.